# Parsimonious Compactly Supported Covariance Models in the Gauss Hypergeometric Class: Identifiability, Reparameterizations, and Asymptotic Properties

### v2

@@ 5 Some special cases and bridges to the Matérn model
If $\kappa=k$ is a nonnegative integer, $d=1$, and $\mu\geq 1$, then
$\begin{split}\mathcal{H}_{k,\mu,a,1}(x)&=\frac{\sqrt{\pi}\,\Gamma\!\left(\frac{\mu}{2}+2k+1\right)}{\Gamma\!\left(k+\tfrac{1}{2}\right)}\left(1-\frac{x}{a}\right)^{\mu+2k}\left(\frac{x}{a}\right)^{k}\\ &\quad\times\sum_{n=0}^{k}\frac{(n+k)!}{2^{2n}n!\,\Gamma\!\left(n+\frac{\mu}{2}+k+1\right)(k-n)!}\left(\frac{(1-\frac{x}{a})^{2}}{\frac{x}{a}}\right)^{n}.\end{split}$
@@
where
$C=\frac{2^{\mu+\frac{d-1}{2}}\Gamma\!\left(\frac{\mu+1}{2}\right)\Gamma\!\left(\frac{\mu+d+1}{2}\right)}{\sqrt{\pi}}.$
Additional closed-form expressions, including spherical models, are reported in Appendix B.
@@
Let $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}{\cal H}_{\kappa,\mu,a,d})$ be two zero-mean Gaussian measures, and let $\kappa>-\frac{1}{2}$ and $\nu=\kappa+\frac{1}{2}$. For any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}{\cal H}_{\kappa,\mu,a,d})$ are equivalent on the paths of $Z(\mathbf{s})$, $\mathbf{s}\in D$, if and only if
$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}}{a^{2\kappa+1}}\,B(\kappa,\mu,d),$ (19)
@@
$B(\kappa,\mu,d)=\frac{2^{2\kappa+1}\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+d+1}{2}+2\kappa\right)}{\Gamma\!\left(\frac{\mu}{2}\right)\Gamma\!\left(\frac{\mu+d}{2}+\kappa\right)}.$ (20)
@@
$a=\alpha\,B(\kappa,\mu,d)^{\frac{1}{1+2\kappa}}:=b(\alpha,\mu,\kappa,d).$
Using this compact-support reparameterization and following the arguments in Bevilacqua et al. (2022), one obtains
@@
It is worth emphasizing that, under this compact-support reparameterization, the role of the shape parameter $\mu$ changes substantially. For fixed smoothness $\kappa$ and scale parameter $\alpha=1$, $\mu$ determines the compact support.
As $\mu\to\infty$, ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ converges to the Matérn model ${\cal MT}_{\kappa+\frac{1}{2},1}$, with maximal integral range given by (13). Conversely, as $\mu$ approaches its lower bound, the compact support decreases, potentially yielding more flexible local dependence structures.
A comparison with the corresponding $\mathcal{GW}_{\kappa,\mu,g(1,\mu,\kappa)}$ model (see (6)) shows that, over a substantial portion of the parameter space, the reparameterized $\mathcal{H}$ models can achieve smaller compact supports. In particular, when $d\geq 2$ or $\kappa\geq 0$,
$b(1,1,\kappa,d)\leq g(1,\mu_{\min}(\kappa,d),\kappa),$
where $\mu_{\min}(\kappa,d)=\kappa+\frac{d+1}{2}$. The case $d=1$ and $\frac{1}{2}<\kappa<0$ is excluded here, as the exact lower bound $\mu_{\min}$ for the $\mathcal{GW}$ model remains an open problem in that regime.
@@ 6 Asymptotic properties of the maximum likelihood estimator of the Hypergeometric model
We study the asymptotic properties of the maximum likelihood (ML) estimator for the parameters of the model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ under both increasing-domain and fixed-domain asymptotics.
Under increasing-domain asymptotics, the parameters $(\sigma^{2},\kappa,\mu,a)$ are consistently estimable, and the ML estimator is asymptotically Gaussian with covariance matrix equal to the inverse Fisher information (Mardia and Marshall, 1984).
Under fixed-domain asymptotics, the situation differs. Using Theorem 1 and the parameterization (11) with $l=\frac{d}{2}+\kappa$, the equivalence condition for two Gaussian measures $P(\sigma_{i}^{2}{\cal H}_{\kappa,\mu_{i},a_{i},d})$, $i=0,1$, sharing the same smoothness parameter, for $d=1,2,3$ and $\min(\mu_{0},\mu_{1})\geq 1+\frac{d}{2}$, is
$\frac{\sigma_{0}^{2}B(\kappa,\mu_{0},d)}{a_{0}^{2\kappa+1}}=\frac{\sigma_{1}^{2}B(\kappa,\mu_{1},d)}{a_{1}^{2\kappa+1}}.$ (22)
Hence only the microergodic parameter
$\frac{\sigma^{2}B(\kappa,\mu,d)}{a^{2\kappa+1}}$
is consistently estimable. If $\mu$ is fixed and known, this reduces to $\sigma^{2}/a^{2\kappa+1}$.
Let $D\subset\mathbb{R}^{d}$ be bounded and $\mathbf{Z}_{n}=(Z(\mathbf{s}_{1}),\ldots,Z(\mathbf{s}_{n}))^{\prime}$ a realization of a zero-mean stationary Gaussian RF with covariance $\sigma^{2}\phi(\|\cdot\|;\boldsymbol{\tau})$, $\phi\in\Phi_{d}$, $\phi(0;\boldsymbol{\tau})=1$. Denote by
$R_{n}(\boldsymbol{\tau})=[\phi(\|\mathbf{s}_{i}-\mathbf{s}_{j}\|;\boldsymbol{\tau})]_{i,j=1}^{n}$
the correlation matrix. The Gaussian log-likelihood is
$\mathcal{L}_{n}(\sigma^{2},\boldsymbol{\tau})=-\frac{1}{2}\Big(n\log(2\pi\sigma^{2})+\log|R_{n}(\boldsymbol{\tau})|+\frac{1}{\sigma^{2}}\mathbf{Z}_{n}^{\prime}R_{n}(\boldsymbol{\tau})^{-1}\mathbf{Z}_{n}\Big).$ (23)
Under the $\mathcal{H}$ model, $\phi(\cdot;\boldsymbol{\tau})=\mathcal{H}_{\kappa,\mu,a,d}$ with $\boldsymbol{\tau}=(\kappa,\mu,a)^{\prime}$. In what follows, $\kappa$ and $\mu$ are assumed fixed and known, so we write $\tau=a$.
Let $\hat{\sigma}_{n}^{2}(a)$ denote the ML estimator of $\sigma^{2}$ for fixed $a>0$:
$\hat{\sigma}_{n}^{2}(a)=\mathbf{Z}_{n}^{\prime}R_{n}(a)^{-1}\mathbf{Z}_{n}/n.$ (24)
The following result establishes strong consistency and asymptotic normality of $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}$. The proof follows the same arguments as Theorem 8 in Bevilacqua et al. (2019).
@@
Let $Z(\mathbf{s})$, $\mathbf{s}\in D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$, $\mu\geq 1+\frac{d}{2}$. Suppose $(\sigma_{0}^{2},a_{0})\in(0,\infty)^{2}$. Then, for fixed $a>0$, as $n\to\infty$,
  1. 1.
    $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\sigma_{0}^{2}/a_{0}^{1+2\kappa}$;
  2. 2.
    $\sqrt{n}\big(\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}-\sigma_{0}^{2}/a_{0}^{1+2\kappa}\big)\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}\big(0,2(\sigma_{0}^{2}/a_{0}^{1+2\kappa})^{2}\big)$.
We next consider joint maximization of (23) with respect to $(\sigma^{2},a)\in(0,\infty)\times I$, $I=[a_{L},a_{U}]$, $0<a_{L}<a_{U}<\infty$. Let
$\hat{\sigma}_{n}^{2}(\hat{a}_{n})=\mathbf{Z}_{n}^{\prime}R_{n}(\hat{a}_{n})^{-1}\mathbf{Z}_{n}/n,\quad\hat{a}_{n}=\arg\max_{a\in I}\mathcal{PL}_{n}(a),$
where $\mathcal{PL}_{n}(a)$ is the profile log-likelihood
$\mathcal{PL}_{n}(a)=-\frac{1}{2}\big(\log(2\pi)+n\log\hat{\sigma}_{n}^{2}(a)+\log|R_{n}(a)|+n\big).$ (25)
The following lemma establishes a monotonicity property of $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}$ as a function of $a$. Its proof is given in Appendix A.
Lemma 1.
Let $S_{n}=\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\}\subset D$. For any $a_{1}<a_{2}$ in $I$,
$\frac{\hat{\sigma}_{n}^{2}(a_{1})}{a_{1}^{1+2\kappa}}\leq\frac{\hat{\sigma}_{n}^{2}(a_{2})}{a_{2}^{1+2\kappa}}\quad\text{if and only if}\quad\mu\geq 4.$
The next theorem establishes the asymptotic properties of $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}$. The proof follows Theorem 9 in Bevilacqua et al. (2019) combined with Lemma 1.
@@
Let $Z(\mathbf{s})$, $\mathbf{s}\in D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$, $\mu\geq 4$. Suppose $(\sigma_{0}^{2},a_{0})\in(0,\infty)\times I$. Then, as $n\to\infty$,
  1. 1.
    $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\sigma_{0}^{2}/a_{0}^{1+2\kappa}$;
  2. 2.
    $\sqrt{n}\big(\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}-\sigma_{0}^{2}/a_{0}^{1+2\kappa}\big)\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}\big(0,2(\sigma_{0}^{2}/a_{0}^{1+2\kappa})^{2}\big)$.
The condition $\mu\geq 4$ ensures the monotonicity property in Lemma 1 and is technical in nature.
Finally, we compare the two reparameterized models $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ and $\sigma^{2}\mathcal{H}_{\kappa,\mu,\alpha\mu,d}$. Under fixed-domain asymptotics and using (22), the associated microergodic parameters are
$\frac{\sigma^{2}}{\alpha^{2\kappa+1}}\quad\text{and}\quad\frac{\sigma^{2}B(\kappa,\mu,d)}{(\alpha\mu)^{2\kappa+1}},$
respectively. In the first case, the microergodic parameter does not depend on $\mu$, so both $\sigma^{2}/\alpha^{2\kappa+1}$ and $\mu$ are consistently estimable. From an inferential perspective, $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ is therefore preferable. For large datasets, however, $\mu$ may be fixed to induce sparsity, in which case the microergodic parameter reduces to $\sigma^{2}/\alpha^{2\kappa+1}$ under both parameterizations.

### v3

@@ 5 Some special cases and bridges to the Matérn model
If $\kappa=k$ is a nonnegative integer, $d=1$, and $\mu\geq 1$, then, for $x\geq 0$,
$\mathcal{H}_{k,\mu,a,1}(x)=\begin{cases}\displaystyle\frac{\sqrt{\pi}\,\Gamma\!\left(\frac{\mu}{2}+2k+1\right)}{\Gamma\!\left(k+\tfrac{1}{2}\right)}\left(1-\frac{x}{a}\right)^{\mu+2k}\left(\frac{x}{a}\right)^{k}\\[6.00006pt] \displaystyle\qquad\times\sum_{n=0}^{k}\frac{(n+k)!}{2^{2n}n!\,\Gamma\!\left(n+\frac{\mu}{2}+k+1\right)(k-n)!}\left(\frac{\left(1-\frac{x}{a}\right)^{2}}{\frac{x}{a}}\right)^{n},&0\leq x<a,\\[11.00008pt] 0,&x\geq a.\end{cases}$
@@
where $C=\frac{2^{\mu+\frac{d-1}{2}}\Gamma\!\left(\frac{\mu+1}{2}\right)\Gamma\!\left(\frac{\mu+d+1}{2}\right)}{\sqrt{\pi}}.$
Additional closed-form expressions are reported in Appendix B, including the spherical models (Chilès and Delfiner, 2012) that are widely used in geostatistical practice and that are recovered as special cases of the proposed hypergeometric family.
@@
Let $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}{\cal H}_{\kappa,\mu,a,d})$ be two zero-mean Gaussian measures, and let $\kappa>-\frac{1}{2}$ , $\nu=\kappa+\frac{1}{2}$ and $\mu>1+d/2$ . For any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}{\cal H}_{\kappa,\mu,a,d})$ are equivalent on the paths of $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$, if and only if
$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}}{a^{2\kappa+1}}\,F(\kappa,\mu,d),$ (19)
@@
$F(\kappa,\mu,d)=\frac{2^{2\kappa+1}\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+d+1}{2}+2\kappa\right)}{\Gamma\!\left(\frac{\mu}{2}\right)\Gamma\!\left(\frac{\mu+d}{2}+\kappa\right)}.$ (20)
@@
$a=\alpha\,F(\kappa,\mu,d)^{\frac{1}{1+2\kappa}}:=b(\alpha,\mu,\kappa,d).$
Using this compact-support reparameterization and following the same arguments in Bevilacqua et al. (2022), one obtains
@@
It is worth emphasizing that, under this compact-support reparameterization, the role of the shape parameter $\mu$ changes substantially. For fixed smoothness $\kappa$ and scale parameter $\alpha$, $\mu$ determines the compact support.
As $\mu\to\infty$, ${\cal H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ converges to the Matérn model ${\cal MT}_{\kappa+\frac{1}{2},\alpha}$, with maximal integral range given by (13).
@@ 6 Asymptotic properties of the maximum likelihood estimator of the Hypergeometric model
We study the asymptotic properties of likelihood-based estimators under fixed-domain (infill) asymptotics for the model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$, with $\kappa$ fixed and known.
Using Theorem 1 and the parameterization (11) with $l=d/2+\kappa$, the equivalence condition for two Gaussian measures $P(\sigma_{i}^{2}\mathcal{H}_{\kappa,\mu_{i},a_{i},d})$, $i=0,1$, sharing the same $\kappa$, for $d=1,2,3$ and $\min(\mu_{0},\mu_{1})>1+d/2$, is
$\frac{\sigma_{0}^{2}\,F(\kappa,\mu_{0},d)}{a_{0}^{2\kappa+1}}=\frac{\sigma_{1}^{2}\,F(\kappa,\mu_{1},d)}{a_{1}^{2\kappa+1}}.$ (22)
Hence, under fixed-domain (infill) asymptotics, the individual parameters $\sigma^{2}$, $\mu$ and $a$ are not consistently identifiable; only the microergodic parameter
$\theta(\sigma^{2},\mu,a):=\frac{\sigma^{2}\,F(\kappa,\mu,d)}{a^{2\kappa+1}}$ (23)
is consistently estimable. In our fixed-domain limit theory we assume $\mu>1+d/2$ for $d=1,2,3$, which is the condition ensuring equivalence of Gaussian measures (see Theorem 1). This requirement is stronger than the mere validity condition for the $\mathcal{H}$ model, which is simply $\mu\geq 1$.
Let $D\subset\mathbb{R}^{d}$ be a bounded domain and let $S_{n}=\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\}\subset D$ be an infill design, i.e. $S_{n}$ becomes dense in $D$ as $n\to\infty$. Let $\mathbf{Z}_{n}=(Z(\mathbf{s}_{1}),\ldots,Z(\mathbf{s}_{n}))^{\top}$ be a realization of a zero-mean stationary Gaussian RF with covariance function $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}(\|\cdot\|)$. Write
$R_{n}(\mu,a):=\bigl[\mathcal{H}_{\kappa,\mu,a,d}(\|\mathbf{s}_{i}-\mathbf{s}_{j}\|)\bigr]_{i,j=1}^{n}$
for the $n\times n$ correlation matrix, and
$\ell_{n}(\sigma^{2},\mu,a):=-\frac{1}{2}\Bigl(n\log(2\pi\sigma^{2})+\log\lvert R_{n}(\mu,a)\rvert+\frac{1}{\sigma^{2}}\mathbf{Z}_{n}^{\top}R_{n}(\mu,a)^{-1}\mathbf{Z}_{n}\Bigr)$ (24)
for the Gaussian log-likelihood. Throughout this section $\kappa$ is fixed and known, while $\mu$ and $a$ may be either fixed or estimated as specified below.
For any $(\mu,a)$, the profile ML estimator of $\sigma^{2}$ is
$\hat{\sigma}_{n}^{2}(\mu,a):=\frac{\mathbf{Z}_{n}^{\top}R_{n}(\mu,a)^{-1}\mathbf{Z}_{n}}{n},$ (25)
and the corresponding plug-in estimator of the microergodic parameter (23) is
$\hat{\theta}_{n}(\mu,a):=\theta\!\bigl(\hat{\sigma}_{n}^{2}(\mu,a),\mu,a\bigr)=\frac{\hat{\sigma}_{n}^{2}(\mu,a)\,F(\kappa,\mu,d)}{a^{2\kappa+1}}.$ (26)
The profile log-likelihood obtained by substituting (25) into (24) is
$\mathcal{PL}_{n}(\mu,a):=-\frac{1}{2}\Bigl(n\log\bigl(2\pi\hat{\sigma}_{n}^{2}(\mu,a)\bigr)+\log\lvert R_{n}(\mu,a)\rvert+n\Bigr).$ (27)
When $\mu$ is fixed and known, the factor $F(\kappa,\mu,d)$ in (23) is a known constant. Thus, the equivalence condition (22) reduces to $\sigma_{0}^{2}/a_{0}^{2\kappa+1}=\sigma_{1}^{2}/a_{1}^{2\kappa+1}$, and the microergodic parameter can equivalently be taken as $\sigma^{2}/a^{2\kappa+1}$ (up to the known multiplicative constant $F(\kappa,\mu,d)$). We keep the notation (23) for uniformity, noting that all statements can be rescaled by $F(\kappa,\mu,d)$ when $\mu$ is fixed.
The following result establishes strong consistency and asymptotic normality of $\hat{\theta}_{n}(\mu,a)$ when both $\mu$ and $a$ are treated as fixed. The proof follows the same arguments as Theorem 8 in Bevilacqua et al. (2019) and is omitted.
@@
Let $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$, $D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with covariance $\sigma_{0}^{2}\,\mathcal{H}_{\kappa,\mu,a_{0},d}$, where $\kappa$ and $\mu>1+d/2$ are fixed and known, and $(\sigma_{0}^{2},a_{0})\in(0,\infty)^{2}$. Then, for any fixed $a>0$, as $n\to\infty$:
  1. (i)
    $\hat{\theta}_{n}(\mu,a)\xrightarrow{a.s.}\theta_{0}:=\sigma_{0}^{2}\,F(\kappa,\mu,d)/a_{0}^{2\kappa+1}$;
  2. (ii)
    $\sqrt{n}\,\bigl(\hat{\theta}_{n}(\mu,a)-\theta_{0}\bigr)\xrightarrow{\mathcal{D}}\mathcal{N}\bigl(0,\,2\theta_{0}^{2}\bigr)$.
We now estimate $a$ by ML over the compact set $I=[a_{L},a_{U}]$:
$\hat{a}_{n}:=\arg\max_{a\in I}\,\mathcal{PL}_{n}(\mu,a),\quad 0<a_{L}<a_{U}<\infty,$ (28)
and study the asymptotic behaviour of the random variable $\hat{\theta}_{n}(\mu,\hat{a}_{n})$. The key technical ingredient is a uniform stochastic control of $\hat{\theta}_{n}(\mu,a)$ over $a\in I$. Lemma 2 (Appendix A) provides this control and implies that replacing a fixed $a$ by the random ML estimator $\hat{a}_{n}$ does not affect the fixed-domain limit distribution of the microergodic estimator.
@@
Under the same assumptions as Theorem 10, and $a_{0}\in I$, let $\hat{a}_{n}$ be as in (28). Then, as $n\to\infty$:
  1. (i)
    $\hat{\theta}_{n}(\mu,\hat{a}_{n})\xrightarrow{p}\theta_{0}$;
  2. (ii)
    $\sqrt{n}\bigl(\hat{\theta}_{n}(\mu,\hat{a}_{n})-\theta_{0}\bigr)\xrightarrow{\mathcal{D}}\mathcal{N}(0,2\theta_{0}^{2})$.
We finally allow $\mu$ to be unknown and estimate $(\mu,a)$ jointly by ML over the compact set $J\times I$, where $J=[\mu_{L},\mu_{U}]$ and $\mu_{L}>1+d/2$:
$(\hat{\mu}_{n},\hat{a}_{n}):=\arg\max_{(\mu,a)\in J\times I}\,\mathcal{PL}_{n}(\mu,a).$ (29)
As in Case 2, the key step is to control $\hat{\theta}_{n}(\mu,a)$ uniformly over the parameter region $J\times I$. Lemma 3 (Appendix A) provides a two-parameter analogue of Lemma 2. Together with Theorem 10, this yields the following result.
Theorem 12.
Under the same assumptions as Theorem 10, with $(\sigma_{0}^{2},\mu_{0},a_{0})\in(0,\infty)\times J\times I$ and $\mu_{0}>1+d/2$, let $(\hat{\mu}_{n},\hat{a}_{n})$ be as in (29). Then, as $n\to\infty$:
  1. (i)
    $\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})\xrightarrow{p}\theta_{0}$;
  2. (ii)
    $\sqrt{n}\,\bigl(\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})-\theta_{0}\bigr)\xrightarrow{\mathcal{D}}\mathcal{N}\bigl(0,\,2\theta_{0}^{2}\bigr)$.
Theorems 1112 show that the fixed-domain limit theory for the microergodic parameter under the hypergeometric model is stable with respect to ML estimation of nuisance covariance parameters over compact sets. Our proofs rely on a different strategy than the classical sandwich arguments based on finite-sample monotonicity (Kaufman and Shaby, 2013; Bevilacqua et al., 2019; Ma and Bhadra, 2022): we establish uniform stochastic control of the plug-in microergodic estimator over the relevant compact parameter sets (see Appendix A). This approach yields more general results than those currently available for related compactly supported families, such as the generalized Wendland model, where the restriction $\mu\geq(d+1)/2+\kappa+3$ is required (Bevilacqua et al., 2019).